Explain how addition and subtraction were used to derive the roman numerals below
1.IV
2.VI
3.XIX
4.LXIV
5.XCIX
1
Expert's answer
2020-05-18T19:45:52-0400

1. "IV=5-1=4"

2. "VI=5+1=6"

3. "XIX=10+(10-1)=10+9=19"

4. "LXIV=50+10+(5-1)=60+4=64"

5. "XCIX=(100-10)+(10-1)=90+9=99"